# Black hole photon ring beyond General Relativity: an integrable parametrization

### v2

@@ Abstract
First, the KOS family is parametrized by two free functions denoted hereafter $(\Delta_{r}(r),\Delta_{y}(y))$ where $y\propto\cos{\theta}$ ) refers to the polar angle. These free functions modify respectively the radial and polar potentials (2.10b) experienced by test particles. Comparing to the previous parametrization [67, 68, 69, 70, 71], the free function $\Delta_{r}(r)$ encompassed a large set of radial deformations treated so far and coincides with the deformation introduced in [71] (denoted $A_{5}(r)$ there). However, to our knowledge, the polar deformation encoded in $\Delta_{y}(y)$ is new. Therefore, the KOS parametrization allows for a new deformation of the Kerr geometry while still preserving its most fundamental hidden symmetries. As we shall see, the new polar deformation allows to parametrize new subtle modifications affecting the critical angles $\theta_{\pm}$ characterizing the polar motion of the test particles which plays a non-trivial role when computing the image of source lensed by a Kerr-like object. Notice that general polar deformations to Kerr have already been considered in [79], but the parametrization introduced in that work does not preserve the rank-2 Killing tensor.
@@
While the functions $\Delta_{r}$ and $\Delta_{y}$ are free, they must satisfy several constraints in order for the spacetime (1.1) to describe our object of interest, i.e. an asymptotically flat black hole. First, asymptotic flatness constrains the asymptotic behavior of $\Delta_{r}$: one must have $\Delta_{r}\sim r^{2}$ as $r\to\infty$. Second, in order for the metric (1.1) to have an event horizon, $\Delta_{r}$ must be such that $\Delta_{r}(r)>0$ for $r>r_{\mathrm{h}}$ and $\Delta_{r}(r_{\mathrm{h}})=0$; $r_{h}$ corresponds to the radius of the outermost horizon. Third, $\Delta_{y}$ must be positive for all possible values of $y$ to conserve the Lorentzian signature. Finally, since the metric exhibits a curvature singularity for $\Sigma=0$, we assume $r_{\mathrm{h}}>0$ in order to avoid the presence of a naked singularity.
@@
This provides the general set of first order equations describing the geodesic motion on an arbitrary representant of the KOS family. The two potentials (2.11a) and (2.11b) govern respectively the radial and polar motion of the test particle [^9] In order to compare (2.11a) and (2.11b) to the standard potentials involved in the radial and polar motion around Kerr potentials, it is useful to write the equation for $\theta(\lambda)$ instead of $y(\lambda)$, using that $\theta=\text{arcos}{(y/a)}$ such that $\left(\frac{\textrm{d}\theta}{\textrm{d}\lambda}\right)^{2}=\frac{1}{a^{2}-y^{2}}\left(\frac{\textrm{d}y}{\textrm{d}\lambda}\right)^{2}\,.$ This leads to the following equation: $\Sigma^{2}\left(\frac{\textrm{d}\theta}{\textrm{d}\lambda}\right)^{2}=\Theta(y)\,,\qquad\Theta(y)=\frac{\Delta_{y}(K-m^{2}y^{2})-(y^{2}E+L)^{2}}{a^{2}-y^{2}}\,.$ One recovers the Kerr potentials using (2.5) and (1.18), i.e. $V_{r}(r)$ $={\mathcal{R}}(r)=\left[(r^{2}+a^{2})E^{\text{BL}}-aL^{\text{BL}}\right]^{2}-\Delta(r^{2}m^{2}+K)\,,$ $V_{y}(y)$ $=\Theta(\theta)=K-\left(\frac{L^{\text{BL}}}{\sin{\theta}}-aE^{\text{BL}}\sin{\theta}\right)^{2}-m^{2}a^{2}\cos^{2}{\theta}\,.$ The classification of the geodesic motion on the Kerr black hole has a long story [95, 96, 97, 48, 98, 99, 100, 101], a full characterization of the set of allowed trajectories being still an active research field [83, 84]. See [102] for a detailed discussion. The different Kerr geodesics can be classified in terms of the range of the constants of motion $(E,L,K,m)$ and the spin $a$ [^10] Notice that it is standard to introduce another constant $Q$ which is well suited to study the equatorial plane, i.e. $y=0$ or $\theta=\pi/2$. In our case, this alternative Carter constant can be written $Q=\frac{V_{y}(0)}{a^{2}}=\frac{K\Delta_{y}(0)-L^{2}}{a^{2}}=\frac{K\Delta_{y}(0)-a^{2}(L^{\text{BL}}-aE^{\text{BL}})^{2}}{a^{2}}$ (2.12) which for Kerr reduces indeed to $Q=K-(L^{\text{BL}}-aE^{\text{BL}})^{2}$ (2.13) . A first condition for causal geodesics is that
@@
In order to capture the different allowed polar motions, it is instructive to recast the polar equation of motion into an equation for a 1d particle in a potential. By setting $\mathrm{d}{\lambda^{\prime}}=\absolutevalue{E}\sqrt{\Delta_{y}}\mathrm{d}{\lambda}/\Sigma$, using the fact that $\Delta_{y}\geq 0$, this equation can be recast into
@@
    >0 or <$\ell>0\mbox{\quad or\quad}\ell<\frac{4\Delta_{y}(0)}{\Delta_{y}^{\prime\prime}(0)}\,.$ (2.31)
@@
  • if $\ell>0$, there are two positions ±$\pm y_{m}$ such that $\Theta(\pm y_{m})=k$ and the polar motion corresponds to oscillations between ±$\pm y_{m}$;
[…]
@@
As seen from the observer, the projected momenta is given by $-\bm{n}$. The relevant information for the photon ring image is carried by the projected momenta of the photon on the observer’s screen. Splitting the vector $-\bm{n}$ as $-\bm{n}=-\bm{n}_{\perp}-\bm{n}_{\parallel}$, one has
$-\bm{n_{\parallel}}=-\frac{p^{y}r_{\mathcal{O}}}{E\sqrt{\Delta_{y}(y_{{\mathcal{O}}})}}\bm{e}_{(y)}-\frac{p^{\varphi}r_{\mathcal{O}}\sqrt{\Delta_{y}(y_{{\mathcal{O}}})}}{E}\bm{e}_{(\varphi)}\,.$ (2.37)
@@
[…]
@@
$\beta$ =±$=\pm\frac{1}{r_{\mathcal{O}}}\left(k-\frac{(y_{\mathcal{O}}^{2}+\ell)^{2}}{\Delta_{y}(y_{{\mathcal{O}}})}\right)^{1/2}\,.$ (2.41b)
[…]
@@
$\beta_{c}$ =±$=\pm\frac{1}{r_{{\mathcal{O}}}\sqrt{\Delta_{y}(y_{{\mathcal{O}}})}}\left[-16r_{0}^{2}\frac{\Delta_{r}(r_{0})^{2}}{\Delta^{\prime}_{r}(r_{0})^{2}}+8r_{0}\frac{\Delta_{r}(r_{0})}{\Delta^{\prime}_{r}(r_{0})^{2}}[2r_{0}\Delta_{y}(y_{\mathcal{O}})+\Sigma(r_{0},y_{\mathcal{O}})\Delta_{r}^{\prime}(r_{0})]-\Sigma^{2}(r_{0},y_{\mathcal{O}})\right]^{1/2}\,.$ (2.42b)
@@
$\vec{r}_{c}={\mathcal{R}}_{c}(r_{0})\vec{u}\qquad\text{with}\qquad\vec{u}=\cos{\phi_{c}(r_{0})}\vec{e}_{\alpha}+\sin{\phi_{c}(r_{0})}\vec{e}_{\beta}\,.$ (2.44)
@@
[…]
@@ Discussion
Analytic characterization of the photon shell: Since the image of the black hole photon ring is generated by the geodesics which are nearly critical, i.e. geodesics which are infinitely close to the unstable critical spherical photon orbits, we have provided a detailed study of this family of geodesics. One of the remarkable outcomes of the KOS family is to allow us to obtain an analytical characterization without fixing the free functions of the metric. Concretely, for an arbitrary KOS geometry, the radial and polar motions are determined by the two potentials (2.18a2.18b) which depend respectively on the radial and polar deformations $\Delta_{r}(r)$ and $\Delta_{y}(y)$. The critical spherical photon orbits which satisfy (2.2.1) are characterized by a reduced angular momentum and Carter constant $(\ell_{c},k_{c})$ given by (2.23). These expressions depend solely on the function $\Delta_{r}(r)$ and the radial equilibrium position $r_{0}$ which can be obtained only once a given choice of $\Delta_{r}(r)$ is made. The extent of the region in which the spherical orbits are allowed, i.e. the photon shell, is constrained by the inequality (2.25). For all the examples studied in this work, the allowed position $r_{0}$ is restricted to a range $[r_{\text{ph}}^{-},r_{\text{ph}}^{+}]$ with $r_{\text{ph}}^{-}>r_{\mathrm{h}}$ where $r_{\mathrm{h}}>0$ stands for the position of the outer horizon. One the other hand, we have provided a detailed classification of the possible polar equilibrium positions for such critical spherical photon orbits. Therefore, for a given KOS representant, the above formula allows one to fully characterize the photon shell geometry.

### v3

@@ Abstract
[…]
@@
While the functions $\Delta_{r}$ and $\Delta_{y}$ are free, they must satisfy several constraints in order for the spacetime (1.1) to describe our object of interest, i.e. an asymptotically flat black hole. First, asymptotic flatness constrains the asymptotic behavior of $\Delta_{r}$: one must have $\Delta_{r}\sim r^{2}$ as $r\to\infty$. Second, in order for the metric (1.1) to have an event horizon, $\Delta_{r}$ must be such that $\Delta_{r}(r)>0$ for $r>r_{\mathrm{h}}$ and $\Delta_{r}(r_{\mathrm{h}})=0$; $r_{h}$ corresponds to the radius of the outermost horizon. Third, $\Delta_{y}$ must be positive for all possible values of $y$ to conserve the Lorentzian signature. Finally, since the metric exhibits a curvature singularity for $\Sigma=0$, we assume $r_{\mathrm{h}}>0$ in order to avoid the presence of a naked singularity. Furthermore, a metric satisfying these constraints might still exhibit a conical or Misner string singularity, depending on the exact shape of the functions $\Delta_{r}$ and $\Delta_{y}$. Requiring these singularities to be removable yields additional constraints on these functions [92, 93] which would need to be further explored.
@@
[…]
@@
[…]
@@
    $\ell>0\mbox{\quad or\quad}\ell<\frac{4\Delta_{y}(0)}{\Delta_{y}^{\prime\prime}(0)}\,.$ (2.32)
@@
  • if $\ell>0$, there are two positions $\pm y_{m}$ such that $\Theta(\pm y_{m})=k$ and the polar motion corresponds to oscillations between $\pm y_{m}$;
  • if $\ifrac{4\Delta_{y}(0)}{\Delta_{y}^{\prime\prime}(0)}<\ell<0$, the movement depends on the value of $k$: either the equation $\Theta(y)=k$ has two solutions $\pm y_{m}$ and the motion corresponds to oscillations around the equatorial plane between these two values, or it has four solutions $\pm y_{m}$ and $\pm y_{v}$ (with $0<y_{v}<y_{m}$) and we can have oscillations either in $\mathopen{[}-y_{m},-y_{v}\mathclose{]}$ and $\mathopen{[}y_{v},y_{m}\mathclose{]}$;
@@
As seen from the observer, the projected momenta is given by $\bm{n}$. The relevant information for the photon ring image is carried by the projected momenta of the photon on the observer’s screen. Splitting the vector $\bm{n}$ as $\bm{n}=-\bm{n}_{\perp}-\bm{n}_{\parallel}$, one has
$\bm{n_{\parallel}}=-\frac{p^{y}r_{\mathcal{O}}}{E\sqrt{\Delta_{y}(y_{{\mathcal{O}}})}}\bm{e}_{(y)}-\frac{p^{\varphi}r_{\mathcal{O}}\sqrt{\Delta_{y}(y_{{\mathcal{O}}})}}{E}\bm{e}_{(\varphi)}\,.$ (2.38)
@@
$\bm{n_{\parallel}}=\alpha\bm{e}_{(\alpha)}+\beta\bm{e}_{(\beta)}\mbox{\quad with\quad}\bm{e}_{(\alpha)}=\bm{e}_{(\varphi)}\mbox{\quad and\quad}\bm{e}_{(\beta)}=-\bm{e}_{(y)}\,$, (2.39)
@@
$\beta$ $=\pm\frac{1}{r_{\mathcal{O}}}\left(k-\frac{(y_{\mathcal{O}}^{2}+\ell)^{2}}{\Delta_{y}(y_{{\mathcal{O}}})}\right)^{1/2}\,.$ (2.42b)
These equations relates the coordinates $(\alpha,\beta)$ of the impact of the photon on the screen with its parameters $(\ell,k)$ and the position and inclination of the observer $(r_{{\mathcal{O}}},y_{{\mathcal{O}}})$. Now, the photon ring is generated by critical photons orbits which have parameters $(\ell,k)$ infinitely close to the critical parameters $(\ell_{c},k_{c})$ characterizing spherical photon orbits [^13] Nonspherical closed photon orbits around the black hole could exist and also contribute to the photon ring. A nonspherical photon orbit exists between a perihelion $r_{\mathrm{p}}$ and aphelion $r_{\mathrm{a}}$ if and only if $\mathcal{V}_{r}(r_{\mathrm{p}})=\mathcal{V}_{r}(r_{\mathrm{a}})=0\mbox{\quad and\quad}\mathcal{V}_{r}(r)\geq 0\mbox{\quad when\quad}r\in\mathopen{[}r_{\mathrm{p}},r_{\mathrm{a}}\mathclose{]}\,.$ For the Kerr metric, this situation in investigated in [108]: due to the roots system of the potential, no such nonspherical orbit is found. For the Kerr off-shell family, despite careful investigation, we found no expression of $\Delta_{r}$ that allows for such orbits. This is the case in particular for all examples studied in this work.. Therefore, these geodesics remain trapped in the region $r_{0}\in\mathopen{[}r_{\mathrm{ph}}^{+},r_{\mathrm{ph}}^{-}\mathclose{]}$ before escaping to infinity and eventually reaching the observer’s screen. The black hole shadow is obtained as the image of these geodesics on a screen far away from the black hole. Therefore, using the expressions (2.23) for the critical parameters $(\ell_{c},k_{c})$ derived in Section 2.2.1, the cartesian coordinates $(\alpha_{c},\beta_{c})$ take the following form:
@@
$\beta_{c}$ $=\pm\frac{1}{r_{{\mathcal{O}}}\sqrt{\Delta_{y}(y_{{\mathcal{O}}})}}\left[-16r_{0}^{2}\frac{\Delta_{r}(r_{0})^{2}}{\Delta^{\prime}_{r}(r_{0})^{2}}+8r_{0}\frac{\Delta_{r}(r_{0})}{\Delta^{\prime}_{r}(r_{0})^{2}}[2r_{0}\Delta_{y}(y_{\mathcal{O}})+\Sigma(r_{0},y_{\mathcal{O}})\Delta_{r}^{\prime}(r_{0})]-\Sigma^{2}(r_{0},y_{\mathcal{O}})\right]^{1/2}\,.$ (2.43b)
@@
[…]
@@
Parameters $M$ $a$ $\ell$ $y_{\mathcal{O}}$ $R_{0}$ $R_{1}$ $R_{2}$ Residuals
Kerr 11 $0.2$ 0 $10.0$ $0.394$ $0.370$ $1.88\text{\times}{10}^{-5}$
SV $1.32$ $0.1$ $0.82$ 0 $10.0$ $0.402$ $0.380$ $1.53\text{\times}{10}^{-5}$
@@ Discussion
Analytic characterization of the photon shell: Since the image of the black hole photon ring is generated by the geodesics which are nearly critical, i.e. geodesics which are infinitely close to the unstable critical spherical photon orbits, we have provided a detailed study of this family of geodesics. One of the remarkable outcomes of the KOS family is to allow us to obtain an analytical characterization without fixing the free functions of the metric. Concretely, for an arbitrary KOS geometry, the radial and polar motions are determined by the two potentials (2.18a2.18b) which depend respectively on the radial and polar deformations $\Delta_{r}(r)$ and $\Delta_{y}(y)$. The critical spherical photon orbits which satisfy (2.2.1) are characterized by a reduced angular momentum and Carter constant $(\ell_{c},k_{c})$ given by (2.23). These expressions depend solely on the function $\Delta_{r}(r)$ and the radial equilibrium position $r_{0}$ which can be obtained only once a given choice of $\Delta_{r}(r)$ is made. The extent of the region in which the spherical orbits are allowed, i.e. the photon shell, is constrained by the inequality (2.25). For all the examples studied in this work, the allowed position $r_{0}$ is restricted to a range $r_{\text{ph}}^{-},r_{\text{ph}}^{+}]$ with $r_{\text{ph}}^{-}>r_{\mathrm{h}}$ where $r_{\mathrm{h}}>0$ stands for the position of the outer horizon. One the other hand, we have provided a detailed classification of the possible polar equilibrium positions for such critical spherical photon orbits. Therefore, for a given KOS representant, the above formula allows one to fully characterize the photon shell geometry.